OpenStudy (anonymous):

the answer for x sqrd -3x-5 for x=-6 49?

OpenStudy (lgbasallote):

please clarify the question...substituting exponents with ^ thingies

OpenStudy (anonymous):

Let me try this again the problem reads: X squared -3X-5 for X=-6

OpenStudy (anonymous):

\[-3x^2-5 \qquad x=-6\]

OpenStudy (phi):

x^2 - 3x -5 with x= -6 (-6)^2 -3(-6) -5 36 + 18 - 5 49 yes. it is 49